# In-Context Learning as Nonparametric Conditional Probability Estimation: Risk Bounds and Optimality

### v1

@@ Abstract
This paper investigates the expected excess risk of In-Context Learning (ICL) for multi-class classification. We model each task as a sequence of labeled prompt samples and a query input, where a pre-trained model estimates the conditional class probabilities of the query. The expected excess risk is defined as the average truncated Kullback–Leibler (KL) divergence between the predicted and ground-truth conditional class distributions, averaged over a specified family of tasks.
We establish a new oracle inequality for the expected excess risk based on KL divergence in multiclass classification. This allows us to derive tight upper and lower bounds for the expected excess risk in transformer-based models, demonstrating that the ICL estimator achieves the minimax optimal rate—up to a logarithmic factor—for conditional probability estimation.
From a technical standpoint, our results introduce a novel method for controlling generalization error using the uniform empirical covering entropy of the log-likelihood function class. Furthermore, we show that multilayer perceptrons (MLPs) can also perform ICL and achieve this optimal rate under specific assumptions, suggesting that transformers may not be the exclusive architecture capable of effective ICL.
Keywords: expected excess risk, in-context learning, multi-class classification, multilayer perceptrons, transformers
@@ 1 Introduction
The transformer architecture (Vaswani et al., 2017) represents a pivotal milestone in modern machine learning. It has been extensively applied across domains including natural language processing (Devlin et al., 2019), computer vision (Dosovitskiy et al., 2021), and reinforcement learning (Chen et al., 2021), catalyzing significant research advancements in the field of large language models (LLMs). Brown et al. (2020) observed that when a model reaches a certain scale, it can emerge the remarkable capability of in-context learning (ICL). Specifically, pre-trained LLMs can perform novel tasks with just a few labeled examples in the prompt, without any parameter updates or fine-tuning. This capability facilitates exceptional few-shot generalization in LLMs while establishing a novel paradigm for few-shot learning (Brown et al., 2020; Garcia et al., 2023). Consequently, it accelerates the deployment of deep learning models, enabling a foundational shift toward LLMs as instruction-following systems.
@@
A recent line of research has quantified ICL by studying transformer models pretrained on linear regression tasks with Gaussian priors (Garg et al., 2022; Akyürek et al., 2023; Li et al., 2023; Raventós et al., 2023; Wu et al., 2024). Concurrently, a growing body of research has revealed a close connection between the forward pass of multi-layer transformer models and multi-step gradient descent algorithms (Akyürek et al., 2023; Von Oswald et al., 2023; Bai et al., 2023; Ahn et al., 2023; Zhang et al., 2024), suggesting that transformers may implement ICL by implicitly simulating gradient descent. Recent empirical findings show that transformers fail to exhibit ICL unless they are pretrained on a sufficiently diverse set of tasks (Raventós et al., 2023), highlighting the crucial role of task diversity in pretraining. The pioneering experimental study by Garg et al. (2022) demonstrated that transformers can in-context learn a variety of fundamental function classes, thereby sparking theoretical analyses of ICL in statistics. Wu et al. (2024) further quantified how the number of pretraining tasks influences the prediction error of ICL in linear regression. Kim et al. (2024) considered deep neural network with one-layer linear transformer structure and established the minimax optimality of the in-context learner from the view of nonparametric regression.
@@
Although substantial research has analyzed the ICL capabilities of large language models, achieving significant progress in understanding this mechanism, there are still notable limitations and challenges in its theoretical analysis.
@@
First, a significant gap exists between the limit assumptions underlying theoretical analyses and the scenarios typically encountered in practical applications. Most existing studies focus on the combined regime of task scaling and context scaling (Abedsoltan et al., 2024), where both the number of pretraining tasks and the context length are assumed to be large. Some analyses further consider the asymptotic limit as both the task scaling and the context scaling approach infinity. However, the practical importance of few-shot scenarios is neglected, as inference is typically constrained by limited context length. For example, GPT-3 was evaluated under few-shot settings  (Brown et al., 2020), where the context length only allowed one or two labeled examples per class in the prompt. To our best knowledge, Wu et al. (2024) rigorously analyzed the impact of task scaling on the prediction risk of ICL at a fixed context length.
@@
Second, there are substantial limitations in both task settings and model architecture, as current theoretical results do not cover multi-classification tasks or complete transformer architecture. Most theoretical studies on ICL concentrate on simplified transformer architectures, such as linear transformers or single-head attention, to facilitate analytical tractability (Zhang et al., 2024; Huang et al., 2024; Kim and Suzuki, 2024; Li et al., 2024a). However, empirical research demonstrates that in a wide range of ICL tasks, multi-head dot-product attention mechanisms offers substantial performance gains over its single-head counterpart (Cui et al., 2024; Xing et al., 2024). For regression problems, some progresses have been made in understanding ICL based on multi-head dot-product attention (Chen et al., 2024a, b; Deora et al., 2024). Li et al. (2024b) and Li et al. (2024c) considered the binary classification and analyzed the training dynamics of single-layer transformer models. Frei and Vardi (2024) investigated in-context classifcation of Gaussian mixtures using a linear transformer model. Meanwhile, multiclass classification, which is fundamental and widely used in deep learning, has not received adequate theoretical attention. Shen et al. (2024) studied the training dynamics of a single-layer linear transformer model for multi-classification of Gaussian mixtures in the context of ICL.
@@
To address these limitations, we focus on multi-classification tasks and develop a theoretical framework for bounding the expected excess risk of ICL under the task-scaling regime. Specifically, we derive non-asymptotic error bounds under the assumption that the number of pretraining tasks is much larger than the context length. We further analyze asymptotic behavior when the number of pretraining tasks tends to infinite while the context length remains fixed. From the perspective of modeling, we frame ICL as a nonparametric conditional probability estimation problem, where the model estimates conditional class probabilities of a query input given labeled prompt examples. We use a deep transformer model with multi-head dot-product attention and analyze its generalization performance in ICL.
@@
    Our work extends existing theoretical frameworks by relaxing main assumptions in experimental and practical settings. Specifically, we relax constraints commonly adopted in previous studies, such as focusing solely on regression task, using simplified transformer architecture, or assuming asymptotic regimes for context scaling. This enables closer alignment with empirical deployments while preserving analytical rigor. Our work bridges critical gaps between theoretical analysis and real-world implementation.
@@
    We establish a new oracle inequality for the expected excess risk based on Kullback–Leibler (KL) divergence in multiclass classification. To the best of our knowledge, this is the first result that decomposes the expected excess risk under KL divergence into approximation, optimization, and generalization errors, where the generalization error is controlled by the uniform empirical covering entropy of the log-likelihood function class.
@@
    Our analysis finds that transformers are not necessarily the uniquely optimal architecture for ICL. By comparing the convergence rates of transformer and MLP architectures, we demonstrate that both can achieve minimax-optimal rates under certain conditions. This reveals that transformers are not the only viable architecture foundation for ICL, which is also corroborated by recent empirical studies (Tong and Pehlevan, 2025; Kratsios and Furuya, 2025).
@@ 2 Related Work
We briefly review related work on ICL and theoretical analyses of transformer architectures across various settings.
@@ 2.1 Theoretical Analyses of Transformers in ICL
Recent theoretical studies have rigorously analyzed various aspects of transformer-based models in ICL. For linear function classes, the statistical complexity and representation power of transformers have been precisely characterized (Zhang et al., 2024; Wu et al., 2024; Cheng et al., 2024), with further extensions to mixture models and broader function classes explored in (Jin et al., 2024; Guo et al., 2024; Collins et al., 2024). In terms of optimization, convergence properties and landscape analyses of transformer-based ICL objectives have been investigated primarily under simplified settings, focusing on linear attention and regression tasks (Zhang et al., 2024; Huang et al., 2024; Kim et al., 2024; Nichani et al., 2024; Kim and Suzuki, 2024). More recently, Chen et al. (2024a) have studied single-layer transformers with multi-head attention in nonlinear regimes, although their analysis remains restricted to shallow architectures and regression. In contrast, we establish expected excess risk bounds for deep transformers with multi-head attention in the multiclass classification setting, thereby substantially extending existing theoretical results.
@@ 2.2 Mechanistic and Algorithmic Perspectives of Transformers in ICL
Recent work has revealed that transformers can emulate classical learning algorithms such as gradient descent and its variants (Akyürek et al., 2023; Von Oswald et al., 2023; Bai et al., 2023; Fu et al., 2024). Studies have also examined global minimizers and critical points that emerge from adaptive gradient descent methods within transformer architectures (Ahn et al., 2023; Mahankali et al., 2024). Additionally, Bayesian analyses suggest that transformers implicitly approximate Bayesian inference, further enhancing the interpretability of their ICL mechanisms (Xie et al., 2022; Müller et al., 2022; Zhang et al., 2022; Panwar et al., 2024; Jeon et al., 2024). Transformers’ ability to perform ICL on data generated by Markov chains has also been investigated (Collins et al., 2024; Edelman et al., 2024; Makkuva et al., 2024), as has their potential for in-context decision-making within reinforcement learning frameworks (Lin et al., 2024; Sinii et al., 2024). Broader theoretical perspectives concerning the generalization behavior, algorithmic interpretations, and meta-learning properties of transformer-based models have been extensively studied by Li et al. (2023), Dai et al. (2023), and Raventós et al. (2023). Analyses of multi-head attention mechanisms, from both representational and Bayesian viewpoints, have further illuminated the learning dynamics in these models (An et al., 2020; Mahdavi et al., 2024). A detailed discussion of these results, however, lies beyond the scope of this paper.
@@ 2.3 ICL Beyond Transformers
Recent literature has begun to investigate whether models beyond transformers can exhibit in-context learning behavior. Tong and Pehlevan (2025) empirically demonstrated that multi-layer perceptrons (MLPs) can achieve ICL performance comparable to transformers on synthetic tasks. Complementing this, Kratsios and Furuya (2025) proved that MLPs possess universal approximation capabilities in an ICL setting. Our theoretical results further support these empirical and theoretical findings by showing that, under comparable conditions, MLPs attain minimax-optimal rates for multiclass classification. This provides evidence that attention-based architectures may not be essential for ICL.
@@ 3 Preliminaries
We write $n]:=\{1,2,\ldots,n\}$. The set of natural numbers including zero is denoted by $\mathbb{N}:=\mathbb{Z}_{\geq 0}$. For a vector $\mathbf{x}=(x_{1},\cdots,x_{p})^{\top}\in\mathbb{R}^{p}$ and a matrix $\mathbf{X}=(X_{ij})\in\mathbb{R}^{p\times N}$, we define the entrywise max-norms $\|\mathbf{x}\|_{\infty}:=\max_{i}|x_{i}|$ and $\|\mathbf{X}\|_{\infty}:=\max_{i,j}|X_{ij}|$, the Euclidean norm $\|\mathbf{x}\|:=(\sum_{i}x_{i}^{2})^{1/2}$, and the Frobenius norm $\|\mathbf{X}\|_{F}:=(\sum_{i,j}X_{ij}^{2})^{1/2}$. The counting norm $\|\mathbf{x}\|_{0}$ denotes the number of nonzero entries in $\mathbf{x}$, and analogously for matrices. For vector-valued functions $\mathbf{f},\mathbf{g}:\Omega\subset\mathbb{R}^{d}\to\mathbb{R}^{K}$, we write $\|\mathbf{f}-\mathbf{g}\|_{\infty}:=\sup_{\mathbf{z}\in\Omega}\max_{j\in[K]}|f_{j}(\mathbf{z})-g_{j}(\mathbf{z})|$, omitting $\Omega$ when clear. For vectors $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{K}$, define $\log(\mathbf{x}/\mathbf{x}^{\prime}):=\log(\mathbf{x})-\log(\mathbf{x}^{\prime})$, applied coordinatewise. For any $\in\mathbb{R}, x\lfloor x\rfloor$ and $\lceil x\rceil$ denote the floor and ceiling, respectively. A standard basis vector in $\mathbb{R}^{K}$ has the form $(0,\ldots,0,1,0,\ldots,0)^{\top}$. The (K1)(K{-}1)-simplex is $\mathcal{S}^{K}:=\{\mathbf{p}\in\mathbb{R}^{K}:\sum_{k}p_{k}=1,\;p_{k}\geq 0\}$. For probability measures $P$ and $Q$, the Kullback–Leibler divergence is $\mathrm{KL}(P\|Q):=\int\log(dP/dQ)\,dP$ if $P\ll Q$, and $\infty$ otherwise. We use the notation $f(x)\lesssim g(x)$ to signify that $f(x)\leq Cg(x)$ for some constant $C>0$. Similarly, $f(x)\gtrsim g(x)$ signifies $f(x)\geq Cg(x)$. The hardmax operator is $\sigma_{H}(\mathbf{x}):=\lim_{c\to\infty}\exp(c\mathbf{x})/\|\exp(c\mathbf{x})\|_{1}$, applied columnwise for matrices. The Hadamard product $\odot$ denotes entrywise multiplication. The metric entropy of the function class $\mathcal{F}$ with respect to the metric $\rho$ is defined as $\mathcal{V}(\delta,\mathcal{F},\rho):=\log\mathcal{N}(\delta,\mathcal{F},\rho)$, where $\mathcal{N}$ is the $\delta$-covering number of $\mathcal{F}$ under the metric $\rho$. Throughout this paper, we use $C$ to denote a generic constant, which may differ from line to line.

### v2

@@ Abstract
This paper investigates the expected excess risk of in-context learning (ICL) for multiclass classification. We formalize each task as a sequence of labeled examples followed by a query input; a pretrained model then estimates the query’s conditional class probabilities. The expected excess risk is defined as the average truncated Kullback-Leibler (KL) divergence between the predicted and true conditional class distributions over a specified family of tasks. We establish a new oracle inequality for this risk, based on KL divergence, in multiclass classification. This yields tight upper and lower bounds for transformer-based models, showing that the ICL estimator achieves the minimax optimal rate (up to logarithmic factors) for conditional probability estimation. From a technical standpoint, our results introduce a novel method for controlling generalization error via uniform empirical entropy. We further demonstrate that multilayer perceptrons (MLPs) can also perform ICL and attain the same optimal rate (up to logarithmic factors) under suitable assumptions, suggesting that effective ICL need not be exclusive to transformer architectures.
Keywords: expected excess risk, in-context learning, multiclass classification, multilayer perceptrons, transformers
@@ 1 Introduction
The transformer architecture (Vaswani et al., 2017) represents a pivotal milestone in modern machine learning. It has been extensively applied across domains including natural language processing (Devlin et al., 2019), computer vision (Dosovitskiy et al., 2021), and reinforcement learning (Chen et al., 2021), catalyzing significant research advancements in the field of large language models (LLMs). Brown et al. (2020) observed that, once scaled sufficiently, LLMs exhibit the remarkable capability of in-context learning (ICL). Specifically, pretrained LLMs can perform novel tasks with just a few labeled examples in the prompt, without any parameter updates or fine-tuning. This capability enables strong few-shot generalization and has established a new paradigm for few-shot learning (Brown et al., 2020; Garcia et al., 2023). Consequently, this paradigm accelerates the deployment of deep learning models, enabling a foundational shift toward LLMs as instruction-following systems.
@@
A recent line of research has quantified ICL by studying transformer models pretrained on linear regression tasks with Gaussian priors (Garg et al., 2022; Akyürek et al., 2023; Li et al., 2023; Raventós et al., 2023; Wu et al., 2024). Concurrently, a growing body of work has revealed a close connection between the forward pass of multi-layer transformer models and multi-step gradient descent algorithms (Akyürek et al., 2023; Von Oswald et al., 2023; Bai et al., 2023; Ahn et al., 2023; Zhang et al., 2024), suggesting that transformers may implement ICL by implicitly simulating gradient descent. Recent empirical findings indicate that transformers fail to exhibit ICL unless they are pretrained on a sufficiently diverse set of tasks (Raventós et al., 2023), underscoring the crucial role of task diversity in pretraining. The pioneering experimental study by Garg et al. (2022) demonstrated that transformer models can in-context learn a variety of fundamental function classes, thereby motivating theoretical analyses of ICL in statistics. Wu et al. (2024) further quantified how the number of pretraining tasks influences the prediction error of ICL in linear regression. Kim et al. (2024) analyzed a model consisting of a deep neural network and a single-layer linear transformer, and established the minimax optimality of the in-context learner from the perspective of nonparametric regression.
@@
Although substantial research has analyzed the ICL capabilities of large language models and achieved significant progress in understanding this mechanism, notable limitations and challenges remain in its theoretical analysis.
@@
First, a significant gap exists between the assumptions of theoretical frameworks and the constraints of practical applications. As noted by Abedsoltan et al. (2024), there are two different aspects of generalization in ICL: task-scaling, where model performance improves as the number of pretraining tasks $T$ increases while keeping the context length $N$ fixed, and context-scaling, where model performance improves as $N$ increases while $T$ remains fixed. Most existing studies focus on the combined regime of task-scaling and context-scaling, where both $T$ and $N$ are assumed to be large. Some analyses further consider the asymptotic limit as both $T$ and $N$ approach infinity. However, the practical importance of few-shot scenarios is neglected, as inference is typically constrained by limited context length. For example, GPT-3 was evaluated under few-shot settings (Brown et al., 2020), where the context length $N$ only allowed one or two labeled examples per class in the prompt. To our knowledge, Wu et al. (2024) provide the first theoretical analysis of the effect of task-scaling on the prediction risk of ICL under a fixed context length.
@@
Second, existing theoretical analyses are constrained by simplified task settings and model architectures. The literature predominantly focuses on linear regression or binary classification, neglecting the widespread practical use of multiclass classification. Architecturally, to maintain analytical tractability, most studies resort to simplified transformer variants, such as linear transformers or single-head attention mechanisms (Zhang et al., 2024; Huang et al., 2024; Kim and Suzuki, 2024; Li et al., 2024a). This stands in stark contrast to empirical evidence, which consistently demonstrates the superior performance of standard multi-head dot-product attention in a wide range of ICL tasks (Cui et al., 2024; Xing et al., 2024). Although some recent work has begun to address multi-head attention for regression problems (Chen et al., 2024a, b; Deora et al., 2024) and binary classification  (Li et al., 2024b, c), the theoretical understanding of multiclass classification remains underdeveloped. Shen et al. (2024) studied multiclass classification of Gaussian mixtures in the context of ICL, but their framework was still limited to single-layer linear transformer models.
@@
To address these limitations, we develop a theoretical framework for multiclass in-context learning that operates under a task-scaling regime with a deep transformer architecture. Specifically, we derive non-asymptotic error bounds on the expected excess risk under the assumption that the number of pretraining tasks $T$ is much larger than the context length $N$. We further analyze the asymptotic behavior as the number of pretraining tasks $T$ tends to infinity while the context length $N$ remains fixed. From a modeling perspective, we frame ICL as a nonparametric conditional probability estimation problem, where the model estimates conditional class probabilities of a query input given labeled prompt examples. To this end, we use a deep transformer model with multi-head dot-product attention and analyze its generalization performance in ICL.
@@
    Our work extends existing theoretical frameworks by relaxing restrictive assumptions that misalign with practical usage. Unlike prior analyses where the context length $N$ is assumed to be large, we focus on few-shot settings which are typically operated in real-world deployments. We also relax other constraints common in previous studies, such as focusing on simplified transformer architectures. These generalizations preserve analytical rigor while enabling closer alignment with real-world deployments, thereby bridging critical gaps between theoretical analysis and practical implementation.
@@
    We establish a new oracle inequality for the expected excess risk based on Kullback–Leibler (KL) divergence in multiclass classification. To the best of our knowledge, this is the first result that decomposes the expected excess risk under KL divergence into approximation, optimization, and generalization errors, with the generalization error controlled by uniform empirical entropy. While prior work such as Schmidt-Hieber (2020) or Bos and Schmidt-Hieber (2022) establishes oracle inequalities for deep neural networks via $L^{\infty}$-metric entropy over the entire covariate space. Our oracle inequality provides a new methodology for risk analysis and can also be applied to other settings of multiclass classification.
@@
    Our risk analysis finds that transformers are not necessarily the uniquely optimal architecture for ICL. By comparing the convergence rates of transformer and MLP architectures, we demonstrate that both can achieve minimax-optimal rates (up to logarithmic factors) under certain conditions. This reveals that transformers may be not the only viable architectural foundation for ICL, which is also corroborated by recent empirical studies (Tong and Pehlevan, 2025; Kratsios and Furuya, 2025).
@@ 2 Related Work
We provide a brief review of related work on ICL and theoretical analyses of transformer architectures in various settings.
@@ 2.1 Theoretical Analyses of Transformers in ICL
Recent theoretical studies have rigorously analyzed various aspects of transformer-based models in ICL. For linear function classes, the statistical complexity and representational power of transformers have been precisely characterized (Zhang et al., 2024; Wu et al., 2024; Cheng et al., 2024), with extensions to mixture models and broader function classes explored in Jin et al. (2024); Guo et al. (2024); Collins et al. (2024). In terms of optimization, convergence properties and landscape analyses of transformer-based ICL objectives have been investigated primarily under simplified settings, focusing on linear attention and regression tasks (Zhang et al., 2024; Huang et al., 2024; Kim et al., 2024; Nichani et al., 2024; Kim and Suzuki, 2024). More recently, Chen et al. (2024a) studied single-layer transformers with multi-head attention in nonlinear regimes, although their analysis is restricted to shallow architectures and regression. In contrast, we establish expected excess risk bounds for deep transformers with multi-head attention in the multiclass classification setting, thereby substantially extending existing theoretical results.
@@ 2.2 Mechanistic and Algorithmic Perspectives of Transformers in ICL
Recent work has shown that transformers can emulate standard learning algorithms such as gradient descent and its variants (Akyürek et al., 2023; Von Oswald et al., 2023; Bai et al., 2023; Fu et al., 2024). Additional studies have examined global minimizers and critical points arising in adaptive gradient descent methods within transformer architectures (Ahn et al., 2023; Mahankali et al., 2024). Bayesian analyses further suggest that transformers implicitly approximate Bayesian inference, offering a probabilistic explanation for ICL mechanisms (Xie et al., 2022; Müller et al., 2022; Zhang et al., 2022; Panwar et al., 2024; Jeon et al., 2024). Transformers’ ability to perform ICL on data generated by Markov chains has also been investigated (Collins et al., 2024; Edelman et al., 2024; Makkuva et al., 2024), as has their potential for in-context decision-making within reinforcement learning frameworks (Lin et al., 2024; Sinii et al., 2024). Broader theoretical perspectives on the generalization behavior, algorithmic interpretations, and meta-learning properties of transformer-based models have been extensively studied by Li et al. (2023), Dai et al. (2023), and Raventós et al. (2023). Analyses of multi-head attention mechanisms, from both representational and Bayesian viewpoints, have further illuminated the learning dynamics in these models (An et al., 2020; Mahdavi et al., 2024). A detailed discussion of these results lies beyond the scope of this paper.
@@ 2.3 ICL Beyond Transformers
Recent literature has begun to investigate whether architectures beyond transformers can exhibit in-context learning behavior. Tong and Pehlevan (2025) empirically demonstrated that multilayer perceptrons (MLPs) can match the ICL performance of transformers on synthetic tasks. Complementing this, Kratsios and Furuya (2025) showed that MLPs are universal approximators in an ICL setting. Our theoretical results further support these findings by demonstrating that, under appropriate conditions, MLPs attain minimax optimal rates for multiclass classification. These results provide evidence that attention-based architectures may not be strictly necessary for ICL.
@@ 3 Preliminaries
We write $n]:=\{1,2,\ldots,n\}$. The set of natural numbers including zero is denoted by $\mathbb{N}:=\mathbb{Z}_{\geq 0}$. For a vector $\mathbf{x}=(x_{1},\cdots,x_{p})^{\top}\in\mathbb{R}^{p}$ and a matrix $\mathbf{X}=(X_{ij})\in\mathbb{R}^{p\times N}$, we define the entrywise max-norms $\|\mathbf{x}\|_{\infty}:=\max_{i}|x_{i}|$, $\|\mathbf{X}\|_{\infty}:=\max_{i,j}|X_{ij}|$, the Euclidean norm $\|\mathbf{x}\|:=(\sum_{i}x_{i}^{2})^{1/2}$, and the Frobenius norm $\|\mathbf{X}\|_{F}:=(\sum_{i,j}X_{ij}^{2})^{1/2}$. The counting norm $\|\mathbf{x}\|_{0}$ denotes the number of nonzero entries in $\mathbf{x}$, and analogously for matrices. For a vector-valued function $\mathbf{f}=(f_{1},\dots,f_{K})^{\top}:\Omega\subset\mathbb{R}^{d}\to\mathbb{R}^{K}$, we define the $L^{\infty}$-norm as $\|\mathbf{f}\|_{\infty}:=\sup_{\mathbf{z}\in\Omega}\max_{j\in[K]}|f_{j}(\mathbf{z})|$, omitting $\Omega$ when clear. For vectors $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{K}$, define $\log(\mathbf{x}/\mathbf{x}^{\prime}):=\log(\mathbf{x})-\log(\mathbf{x}^{\prime})$, applied coordinatewise. For any $\in\mathbb{R}, x\lfloor x\rfloor$ and $\lceil x\rceil$ denote the floor and ceiling, respectively. A standard basis vector in $\mathbb{R}^{K}$ has the form $(0,\ldots,0,1,0,\ldots,0)^{\top}$. The (K1)(K-1)-simplex is $\mathcal{S}^{K}:=\{\mathbf{p}=(p_{1},\dots,p_{K})^{\top}\in\mathbb{R}^{K}:\sum_{k}p_{k}=1,\;p_{k}\geq 0\}$. For probability measures $P$ and $Q$, the Kullback–Leibler divergence is $\mathrm{KL}(P\|Q):=\int\log(dP/dQ)\,dP$ if $P\ll Q$, and $\infty$ otherwise. We use the notation $f(x)\lesssim g(x)$ to signify that $f(x)\leq Cg(x)$ for all $x$ and some constant $C>0$. Similarly, $f(x)\gtrsim g(x)$ signifies $f(x)\geq Cg(x)$. For $\mathbf{x}\in\mathbb{R}^{p}$, we write $\|\mathbf{x}\|_{1}=\sum_{i}|x_{i}|$. The hardmax operator is $\sigma_{H}(\mathbf{x}):=\lim_{c\to\infty}\exp(c\mathbf{x})/\|\exp(c\mathbf{x})\|_{1}$, applied columnwise for matrices. The Hadamard product $\odot$ denotes entrywise multiplication. The metric entropy of the function class $\mathcal{F}$ with respect to the metric $\rho$ is defined as $\mathcal{V}(\delta,\mathcal{F},\rho):=\log\mathcal{N}(\delta,\mathcal{F},\rho)$, where $\mathcal{N}(\delta,\mathcal{F},\rho)$ is the $\delta$-covering number of $\mathcal{F}$ under the metric $\rho$. Throughout this paper, we use $C$ to denote a generic constant, which may differ across appearances.